How many liters of pure water should be mixed with a 13-L solution of 90% acid to produce a mixture that is 70% water?
Answer by Alan3354(69443)

How many liters of pure water should be mixed with a 13-L solution of 90% acid to produce a mixture that is 70% water?
----------------
70% water = 30% acid
---------
13 liters of 90% acid has 12*0.9 = 11.7 liters of acid
If 11.7 l is 30%, then the total solution = 11.7/0.3 = 39 liters
39 - 13 = 26 liters of water to add